# https://zielon.github.io/gem/

### v1

@@ 5.3 GEM Compression Ablation
5.4 Universality of GEM
We designed GEM to be a universal method capable of distilling any 3DGS-based avatar solution [28, 38, 74, 15, 53] into a lightweight representation, as long as normalized input across training frames is provided. Our results show that a compact representation of the linear basis produces state-of-the-art results in terms of quality and speed. Figure 11 demonstrates how we can apply GEM to different primitive generators (AG [38], GA [53], Our Net). Note that to achieve wrinkle-level details, the generator itself has to produce high-quality output. As can be seen, we can apply our distillation technique to existing methods, making them light and compact. Moreover, GEM is well-suited for commodity devices, as it generates a new set of Gaussian primitives with only a single dot product needed between coefficients and the basis. We believe that this potential has great implications for downstream tasks such as holoportation, audio-driven avatars, or virtual reality.
@@ 6 Human Head Avatar Compression
7 Discussion
Despite state-of-the-art results for human face avatars, our method has still certain limitations which we would like to discuss. One of the major drawbacks of the PCA-based models is their global extent. Which ultimately means that we cannot control local changes and produce more combinations of local features beyond the standard model. Thus, further work could include incorporating a localized PCA basis [44] for better avatar control, which could potentially enable a wider range of expressions outside the training set. Another limitation is the personalization, for new subjects a new representation has to be learned from multi-view data. An interesting future avenue would be to create a statistical model across different subjects.
8 Conclusion
We have proposed Gaussian Eigen Model, a linear appearance model that can represent photo-realistic facial avatars. The simplicity of this appearance model results in massively reduced compute requirements in comparison to CNN-based avatar methods. Moreover, the lightweight representation could improve the management, sharing, and applicability of avatars. The universal distillation approach can be applied to many existing methods, making them available for compression. We demonstrate how such a representation can be used in different scenarios like self-reenactment and cross-person animation. We also show that real-time applications are possible. We believe that the proposed GEM is a stepping stone toward lightweight avatar animation on consumer-level hardware.
Gaussian Eigen Models for Human Heads
–Supplemental Materials– Wojciech Zielonka1 Timo Bolkart2 Thabo Beeler2 Justus Thies1,3
Ground Truth
$128^{2}$
$256^{2}$
$512^{2}$
#10#30#50
9 Compression Ablation Study
Expanding upon this foundation, GEM proposes a novel technique involving an ensemble of eigenbasis of 3D Gaussian attributes for achieving a photorealistic human head appearance. This representation exhibits significant adaptability concerning both quality and size, leveraging a fundamental trait of linear basis that proves beneficial when applied to diverse devices with varying capabilities in digital human applications. Figures 12 and 13 illustrate the qualitative and quantitative results of GEM. Notably, even under substantial compression (utilizing only ten principal components), our approach consistently yields high-quality outcomes. More examples can be found in Figures (17, 15, 18, 16)
$128^{2}$
$256^{2}$
$512^{2}$
#10#30#50
10 Additional Dataset Evaluation
In Figure 19, we provide further evaluation of our approach using the Multiface dataset [72]. This dataset encompasses short sequences of facial expressions, ranging from "relaxed mouth open" to "show all teeth" or "jaw open huge smile." The expressions vary widely in length and complexity, presenting a considerable challenge for analysis. It’s important to note that this dataset does not provide a parametric 3DMM; instead, it offers meshes in full correspondence. However, as mentioned in the main text, our method remains adaptable in this context. By leveraging the deformation gradient [60] to transform points from canonical space into deformed space, and assuming consistent UV parametrization of input meshes, we can successfully navigate between these spaces. As depicted in Figure 19, our network demonstrates the ability to extrapolate to novel expressions, even amidst highly challenging facial poses.
11 Broader Impact
@@
12 Future Applications & Discussion
Moreover, GEM neural network uses meshes to obtain normal maps as input to the Gaussian map regressor. However, meshes are limited by resolution and expressiveness, one way of improving on that would be to use NPHM by [13] by Giebenhain et al. and the follow-up work [61, 14, 31] to further increase the expressiveness of the model by explicitly capturing regions like hair or teeth.
@@ 12 Future Applications & Discussion
Figure 19: The Multiface dataset, introduced by Wu et al. [72], comprises actors performing scripted expressions in short segments. A notable challenge arises due to the occurrence of several expressions, like "show all teeth," appearing only once in the dataset. This poses a difficulty during testing, particularly when the network is required to extrapolate. Here we showcase the outcomes of the test sequences to illustrate the effectiveness of our CNN-based network in capturing diverse and challenging facial poses, demonstrating its robustness despite the inherent complexity of the dataset.

### v2

@@ B.2 CNN Training Details
Our method and all the baselines were trained using the same multiview input data sourced from the dataset provided by Qian [57], which includes multiview images from the NeRSamble dataset [33] as well as tracked meshes.
Appendix C Compression Ablation Study
Expanding upon this foundation, GEM proposes a novel technique involving an ensemble of eigenbases of 3D Gaussian attributes for achieving a photorealistic human head appearance. This representation exhibits significant adaptability concerning both quality and size, leveraging a fundamental trait of linear basis that proves beneficial when applied to diverse devices with varying capabilities in digital human applications. Figures 17 and 18 illustrate the qualitative and quantitative results of GEM. Notably, even under substantial compression (utilizing only ten principal components), our approach consistently yields high-quality outcomes. More examples can be found in Figures 19, 20, 21 and 22.
Appendix D Human Head Avatar Compression
#Comp $128^{2}$ $256^{2}$ $512^{2}$
10 7 28 113
30 20 83 333
50 34 138 553
Ours Net 82 109 178
GA StyleUnet 487 529 636
Table 6: Memory consumption (in MB with float32) of GEM depends on the texture resolution and number of components. Our model shows much better granularity compared to the fixed size of neural networks and can be adjusted on the fly depending on the hardware.
#Comp $128^{2}$ $256^{2}$ $512^{2}$
10 31.47 31.90 31.80
30 33.46 34.30 34.26
50 33.79 34.75 34.73
Table 7: PSNR color error in dB for one actor with a different number of principle components and Gaussian map resolutions. Despite heavy compression (10 principal components), the avatar is still of high quality. More details are in the supplementary material.
@@ Appendix D Human Head Avatar Compression
Appendix E Additional Dataset Evaluation
In Figure 23, we provide further evaluation of our approach using the Multiface dataset [77]. This dataset encompasses short sequences of facial expressions, ranging from ”relaxed mouth open” to ”show all teeth” or ”jaw open huge smile.” The expressions vary widely in length and complexity, presenting a considerable challenge for analysis. It’s important to note that this dataset does not provide a parametric 3DMM; instead, it offers meshes in full correspondence. However, as mentioned in the main text, our method remains adaptable in this context. By leveraging the deformation gradient [64] to transform points from canonical space into deformed space, and assuming consistent UV parametrization of input meshes, we can successfully navigate between these spaces. As depicted in Figure 23, our network demonstrates the ability to extrapolate to novel expressions, even amidst highly challenging facial poses.
Appendix F Broader Impact
@@
Appendix G Future Applications & Discussion
Moreover, our neural network based appearance model uses meshes to obtain normal maps as input to the Gaussian map regressor (similar to the baselines). However, meshes are limited by resolution and expressiveness, one way of improving on that would be to use NPHM by Giebenhain et al. [14] and the follow-up work [65, 15, 32] to further increase the expressiveness of the model by explicitly capturing regions like hair or teeth.
@@ Appendix G Future Applications & Discussion
Refer to captionGround TruthOurs NetGA. [57]AG. [41]INSTA [88]
@@
Ground Truth
$128^{2}$
$256^{2}$
$512^{2}$
#10#30#50
$128^{2}$
$256^{2}$
$512^{2}$
#10#30#50
Figure 22: Compression error in PSNR (db) depending on the number of bases (10, 30, 50) and resolution of the Gaussian maps ($128^{2}$, $256^{2}$, $512^{2}$).
@@
Figure 23: The Multiface dataset, introduced by Wu et al. [77], comprises actors performing scripted expressions in short segments. A notable challenge arises due to the occurrence of several expressions, like ”show all teeth,” appearing only once in the dataset. This poses a difficulty during testing, particularly when the network is required to extrapolate. Here we showcase the outcomes of the test sequences to illustrate the effectiveness of our CNN-based network in capturing diverse and challenging facial poses, demonstrating its robustness despite the inherent complexity of the dataset.